# ARM: Advantage Reward Modeling for Long-Horizon Manipulation

### 03 Apr 2026 v1

@@ Abstract
Project page: https://aiming1998.github.io/ARM/
@@ 2.2 Reward-Aligned Behavior Cloning (RA-BC)
However, traditional weighting paradigms are limited by a critical bottleneck: they inherently rely on explicit environment rewards to fit global value functions, which are notoriously inaccessible in vision-based, real-world settings. To bypass this, recent methods like SARM [6] introduced the Reward-Aligned Behavior Cloning (RA-BC) framework, leveraging a stage-aware reward model instead of environment rewards. While effective at mitigating data quality issues, SARM trades the reward bottleneck for a new constraint: it heavily relies on prohibitive manual language annotations. In contrast, our proposed ARM eliminates the need for both explicit reward engineering by extracting advantage signals purely through relative progress comparisons.
@@ 3.2 Advantage Reward Modeling
The Advantage Reward Model (ARM) is designed to resolve the perceptual ambiguities inherent in isolated frames by shifting the reward estimation paradigm from absolute progress regression to relative advantage classification. Unlike traditional ”Multi-Input Single-Output” (MISO) models [6, 21] that collapse temporal context into a single scalar, ARM formulates reward estimation as a Multi-Input Multi-Output (MIMO) sequence learning problem (Fig. 2). This design allows the model to contextualize local observations within a short-term history, analogous to how humans review recent temporal context to disambiguate intent and action quality.
@@ 3.2.1 MIMO Transformer Architecture
Dual-Head Learning Objective
@@
3.2.2 Lightweight Tri-state Auto Labeling Strategy.
@@ 3.4.1 Length-adaptive Gain Formulation
where $P_{t}$ denotes the progress value by global progress reconstruction, $L_{\text{seq}}$ represents the total length of the current episode, and $\bar{L}$ is the average episode length across the entire dataset. This normalization ensures that the derived advantage reflects the relative efficiency of a specific action sequence, effectively decoupling the reward signal from the absolute duration of the task.
@@ 4.1 Experimental Setup
Data was collected using an AgileX ALOHA[9] bimanual teleoperation system with randomized table heights for enhanced generalization. Detailed implementation details are in Section B.
@@
We evaluate three annotation paradigms: (i) VLM-based Labeling using Qwen3-VL [29] for temporal grounding of subtask boundaries; (ii) Manual Subtask Segmentation by human experts; and (iii) our proposed Tri-state Labeling.
@@ 4.2 Reward Model Performance
Table 1: Quantitative Evaluation of Reward Models. All models are evaluated on a validation set of 50 trajectories. “MSE” measures the trajectory reconstruction fidelity against Ground Truth (GT) progress (normalized to $[0,1]$). The bottom section reports the Success Identification Accuracy, assessing the Completion Head’s ability to correctly classify the final state of Standard (SE, 12 successful episodes), and Failure (FE, 12 failed episodes) trajectories. Best performances are highlighted in bold.
@@
Figure 5: Qualitative comparison of progress reconstruction. We visualize the progress curves of SARM and ARM against the Ground Truth (GT) for a representative episode. While SARM struggles with non-monotonic behaviors, ARM reconstructs a smooth, high-fidelity curve that closely tracks the GT, even during regressive adjustments.
@@ 4.3 Efficiency and Quality of Reward Labeling
Table 3: Labeling Efficiency Comparison. We compare the sample throughput per 8-hour shift across different protocols.
@@
As shown in Table 3, our tri-state protocol achieves significant efficiency gains. By simplifying annotation from precise temporal boundary localization to discrete classification, human annotators achieve 250 samples per 8-hour shift—a 2.5× speedup over the baseline (100 samples). This simplified formulation enables massive scaling: our Auto Tri-state pipeline processes >2,000>2,000 samples per 8 hours, achieving a $20\times$ speedup over human baselines.
@@ 4.5 Policy Performance Analysis
  • (2) RA-BC (GR00T+SARM): A Reward-Aware BC policy [6] re-weighted by the absolute progress signals from the SARM model.
@@
As summarized in Table 2, the standard BC-Baseline suffers from a suboptimal success rate (62.1%) and lower operational efficiency. This is primarily due to the multi-modal noise and “sluggish” trajectories inherent in the full dataset, which typical BC fails to filter or prioritize. While RA-BC (SARM) improves the success rate to 78.5% through subtask-based weighting, it remains constrained by the lack of fine-grained advantage estimation for error-correction behaviors.
@@ 5 Conclusion
We propose Advantage Reward Modeling (ARM), a framework that addresses the reward engineering bottleneck in long-horizon robotic manipulation tasks. By modeling relative advantages, ARM overcomes inconsistency and high costs of traditional dense labeling. We introduce a tri-state labeling strategy that reduces cognitive load for annotators while providing high-fidelity supervision signals and enabling automated labeling. In a challenging towel-folding task, ARM with Advantage-Weighted Behavior Cloning achieves a $99.4\%$ success rate, outperforming existing Vision-Language-Action baselines. ARM provides a scalable and robust solution for training high-performance policies.
@@ Appendix A VLM Prompting Details
For VLM-based annotation, following SARM [6], we use Qwen3-VL [29] from LeRobot [4] to segment demonstration videos into temporally ordered subtask intervals using a fixed vocabulary.
@@ Appendix B Implementation Details
ARM utilizes a pre-trained CLIP ViT-B/32 [30] as the vision-text encoder, followed by a Transformer-based Sequential Aggregator with a causal 5-frame window (sampled at 1Hz). The joint objective is defined as $\mathcal{L}_{\text{ARM}}=\lambda_{\text{int}}\mathcal{L}_{\text{int}}+\lambda_{\text{succ}}\mathcal{L}_{\text{succ}}$, where we employ Focal Loss for task completion and cross-entropy for tri-state interval classification. Complete hyperparameters are summarized in Table 6.
@@
Based on the GR00T-N1.5 [1] VLA foundation, our policy uses Advantage-Weighted Behavior Cloning where sample weights $w$ are derived from ARM-predicted gains $\Delta G_{t}$. Training configurations are detailed in Table 7.
@@ Appendix C ARM Inference Results
Figure 7: Visualization of ARM Inference Results. The left panels show the third-person view of the bimanual towel-folding task at $t=69s$ and $t=70s$. The right panels display the corresponding progress curves: predicted progress $P_{pred}$ (blue) and ground truth $P_{gt}$ (green). ARM accurately captures the non-monotonic progress “dip” caused by a regressive adjustment, with the interval head correctly outputting $\Delta_{pred}=-1$.
@@
As captured in the transition from $t=69$s ($P_{\text{pred}}=86.15\%$) to $t=70$s ($P_{\text{pred}}=84.62\%$), ARM’s interval head successfully identifies this trend, consistently predicting regressive signals ($\Delta_{pred}=-1$, as shown in the status text). This causes the reconstructed progress curve (blue line) to exhibit a precise downward “dip” that closely aligns with the ground truth (green line).
@@ Appendix D Real-World Implementation Details
The real-world data collection and policy deployments were conducted using an AgileX master-slave teleoperation system (illustrated in Fig. 8). The hardware platform utilizes a 6-Degree-of-Freedom (6-DoF) bimanual robot configuration.
@@ Appendix E Folding Precision Evaluation Protocol
We define a quantitative folding precision score ranging from 0 to 5 to evaluate the quality of towel folding results: 5 points: The folding task is fully completed, with a folding precision within 1 cm. 4 points: The folding task is fully completed, with a folding precision between 1 cm and 2 cm. 3 points: The folding task is fully completed, with a folding precision between 2 cm and 3 cm. 2 points: The towel is successfully flattened, but the final folding is not completed, though partial folding steps are finished. 1 point: The towel is successfully flattened, but no valid folding steps are performed. 0 points: No task steps are successfully completed.

### 21 Apr 2026 v2

@@ Abstract
Project page: https://aiming1998.github.io/ARM
@@ 2.2 Reward-Aligned Behavior Cloning (RA-BC)
However, traditional weighting paradigms are limited by a critical bottleneck: they inherently rely on explicit environment rewards to fit global value functions, which are notoriously inaccessible in vision-based, real-world settings. To bypass this, recent methods like SARM [6] introduced the Reward-Aligned Behavior Cloning (RA-BC) framework, leveraging a stage-aware reward model instead of environment rewards. While effective at mitigating data quality issues, SARM trades the reward bottleneck for a new constraint: it heavily relies on prohibitive manual language annotations. In contrast, our proposed ARM eliminates the need for explicit reward engineering by extracting advantage signals purely through relative progress comparisons.
@@ 3.2 Advantage Reward Modeling
The Advantage Reward Model (ARM) is designed to resolve the perceptual ambiguities inherent in isolated frames by shifting the reward estimation paradigm from absolute progress regression to relative advantage classification. Unlike traditional “Multi-Input Single-Output” (MISO) models [6, 21] that collapse temporal context into a single scalar, ARM formulates reward estimation as a Multi-Input Multi-Output (MIMO) sequence learning problem (Fig. 2). This design allows the model to contextualize local observations within a short-term history, analogous to how humans review recent temporal context to disambiguate intent and action quality.
@@ 3.2.1 MIMO Transformer Architecture
Dual-Head Learning Objective.
@@
3.2.2 Lightweight Tri-state Auto Labeling Strategy
@@ 3.4.1 Length-adaptive Gain Formulation
where $P_{t}$ denotes the progress value obtained via global progress reconstruction, $L_{\text{seq}}$ represents the total length of the current episode, and $\bar{L}$ is the average episode length across the entire dataset. This normalization ensures that the derived advantage reflects the relative efficiency of a specific action sequence, effectively decoupling the reward signal from the absolute duration of the task.
@@ 4.1 Experimental Setup
Data was collected using an AgileX ALOHA [9] bimanual teleoperation system with randomized table heights for enhanced generalization. Detailed implementation details are in the Supplementary Materials.
@@
We evaluate three annotation paradigms: (i) VLM-based Labeling implemented in LeRobot [4], using Qwen3-VL [29] for temporal grounding of subtask boundaries; (ii) Manual Subtask Segmentation by human experts; and (iii) our proposed Tri-state Labeling.
@@ 4.2 Reward Model Performance
Table 1: Quantitative Evaluation of Reward Models. All models are evaluated on a validation set of 50 trajectories. “MSE” measures the trajectory reconstruction fidelity against GT progress (normalized to $[0,1]$). The bottom section reports the Success Identification Accuracy, assessing the Completion Head’s ability to correctly classify the final state of Standard (SE, 12 successful episodes), and Failure (FE, 12 failed episodes) trajectories. Best performances are highlighted in bold.
@@
Figure 5: Qualitative comparison of progress reconstruction. We visualize the progress curves of SARM and ARM against the GT for a representative episode. While SARM struggles with non-monotonic behaviors, ARM reconstructs a smooth, high-fidelity curve that closely tracks the GT, even during regressive adjustments.
@@ 4.3 Efficiency and Quality of Reward Labeling
Table 3: Labeling Efficiency Comparison. Annotation throughput comparison between human and automated labeling protocols per 8-hour shift.
@@
As shown in Table 3, our tri-state protocol achieves significant efficiency gains. By simplifying annotation from precise temporal boundary localization to discrete classification, human annotators achieve 250 samples per 8-hour shift—a 2.5× speedup over the baseline (100 samples). This simplified formulation enables massive scaling: our Auto Tri-state pipeline processes >400,000>400,000 samples per 8 hours, achieving a $133\times$ speedup over human baselines.
@@ 4.5 Policy Performance Analysis
  • (2) RA-BC (GR00T+SARM): Reward-Aligned Behavior Cloning [6] re-weighted by SARM progress signals.
@@
As summarized in Table 2, the Baseline suffers from a suboptimal success rate (62.1%) and lower operational efficiency. This is primarily due to the multi-modal noise and “sluggish” trajectories inherent in the full dataset, which typical BC fails to filter or prioritize. While RA-BC(GR00T+SARM) improves the success rate to 78.5% through subtask-based weighting, it remains constrained by the lack of fine-grained advantage estimation for error-correction behaviors.
@@ 5 Conclusion
We propose Advantage Reward Model (ARM), a framework that addresses the reward engineering bottleneck in long-horizon robotic manipulation tasks. By modeling relative advantages, ARM overcomes inconsistency and high costs of traditional dense labeling. We introduce a tri-state labeling strategy that reduces cognitive load for annotators while providing high-fidelity supervision signals and enabling automated labeling. In a challenging towel-folding task, ARM with Advantage-Weighted Behavior Cloning achieves a $99.4\%$ success rate, outperforming existing Vision-Language-Action baselines. ARM provides a scalable and robust solution for training high-performance policies.
@@ Appendix B Implementation Details
ARM utilizes a pre-trained CLIP ViT-B/32 as the vision-text encoder, followed by a Transformer-based Sequential Aggregator with a causal 5-frame window (sampled at 1Hz). The joint objective is defined as $\mathcal{L}_{\text{ARM}}=\lambda_{\text{int}}\mathcal{L}_{\text{int}}+\lambda_{\text{succ}}\mathcal{L}_{\text{succ}}$, where we employ Focal Loss for task completion and cross-entropy for tri-state interval classification. Complete hyperparameters are summarized in Table 6.
@@
Based on the GR00T-N1.5 VLA foundation, our policy uses Advantage-Weighted Behavior Cloning where sample weights $w$ are derived from ARM-predicted gains $\Delta G_{t}$. Training configurations are detailed in Table 7.
@@ Appendix C ARM Inference Results
Figure 7: Visualization of ARM Inference Results. The left panels show the third-person view of the bimanual towel-folding task at $t=69s$ and $t=70s$. The right panels display the corresponding progress curves: predicted progress $P_{pred}$ (blue) and ground truth $P_{gt}$ (green). ARM accurately captures the non-monotonic progress “dip” caused by a regressive adjustment, with the Multi-frame Advantage head correctly outputting $\Delta_{\textbf{pred}}=-1$.
@@
As captured in the transition from $t=69$s ($P_{\text{pred}}=86.15\%$) to $t=70$s ($P_{\text{pred}}=84.62\%$), the Multi-frame Advantage head successfully identifies this trend, consistently predicting regressive signals ($\Delta_{\textbf{pred}}=-1$, as shown in the status text). This causes the reconstructed progress curve (blue line) to exhibit a precise downward “dip” that closely aligns with the ground truth (green line).
@@ Appendix D Real-World Implementation Details
The real-world data collection and policy deployment were conducted using an AgileX master-slave teleoperation system (illustrated in Fig. 8). The hardware platform utilizes a 6-Degree-of-Freedom (6-DoF) bimanual robot configuration.
@@ Appendix E Folding Precision Evaluation Protocol
We define a quantitative folding precision score ranging from 0 to 5 to evaluate the quality of towel folding results:
  • 5 points: The folding task is fully completed, with a folding precision within 1 cm.
  • 4 points: The folding task is fully completed, with a folding precision between 1 cm and 2 cm.
  • 3 points: The folding task is fully completed, with a folding precision between 2 cm and 3 cm.
  • 2 points: The towel is successfully flattened, but the final folding is not completed, though partial folding steps are finished.
  • 1 point: The towel is successfully flattened, but no valid folding steps are performed.
  • 0 points: No task steps are successfully completed.